Pt-200 undergoes beta decay. What is the product nucleus? Enter your answer using the same format, i.e, symbol-mass number

In-120 undergoes beta decay. What is the product nucleus? Enter your answer using the same format, i.e, symbol-mass number
Answer #1

1)

Let the product be X

Balance mass number on both sides

mass of Pt = mass of e + mass of X

1*200 = 1*0 + M

M = 200

Balance atomic number on both sides

atomic of Pt = atomic of e + atomic of X

1*78 = 1*-1 + A

A = 79

Atomic number 79 is for element Au

So, the product is

200-Au

Answer: 200-Au

2)

Let the product be X

Balance mass number on both sides

mass of In = mass of e + mass of X

1*120 = 1*0 + M

M = 120

Balance atomic number on both sides

atomic of In = atomic of e + atomic of X

1*49 = 1*-1 + A

A = 50

Atomic number 50 is for element Sn

So, the product is

120-Sn

Answer: 120-Sn